OpenStudy (anonymous):

Which of the following metric prefixes is the smallest A) Giga B)Deka C)Mega D)Kilo

OpenStudy (anonymous):

I think kilo

OpenStudy (anonymous):

its not kilo

OpenStudy (anonymous):

Deka =10^1 kilo= 10^3 giga=10^9 mega= 10^6

OpenStudy (anonymous):

deka is 10? oh